OpenStudy (anonymous):

Which of the following sets of vectors are vector spaces? Justify a) R^2 U R^3 b) {|x belongs to R} c) {|k is an integer} d) {|x belongs to R} U {<0,y>|y belongs to R} f) Span({<1,3,2>,<-1,0,3>})

OpenStudy (anonymous):

e isn't a vector space. (9,0) is in the space, and (0,7) is in the space, but (9,0)+(0,7)=(9,7) is not in the space.

OpenStudy (anonymous):

c is only a vector space if the scalar field is the integers. If it is the Reals, then c is not a vector space.

OpenStudy (anonymous):

ok...how about a? i think it's not a vector space. am i right?

OpenStudy (anonymous):

a is the union of R2 and R3?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So some elements will be of form \[\left(\begin{matrix}a \\ b\end{matrix}\right)\]and others of form \[\left(\begin{matrix}a \\ b \\ c\end{matrix}\right)\]?

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

Hmmm... I think it would be fine, just what is the way to add a vector from R2 and a vector from R3?

OpenStudy (anonymous):

Not too sure...Is it even possible to add <a,b> with <c,d,e>?

OpenStudy (anonymous):

and according to your answer for (c), (b) wont be a vector space. why so?

OpenStudy (anonymous):

Not normally just like that. But there are ways around it. For example, \[\mathbb{R}^{2}\subseteq \mathbb{R}^{3}\], and you could just write all of the elements from R2 as \[\left(\begin{matrix}a \\ b \\ 0\end{matrix}\right)\]

OpenStudy (anonymous):

so, what you're saying is R2 U R3 is essentially R3 and hence a vector space. right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

But b is a vector space. The reason c isn't is that it requires that all vectors be of the form (k, 2k) where k is an integer. But you could multiply a vector by say 7.59378, and then you couldn't get a vector of form (k,2k) with k being an integer. The one in b has no such restriction.

OpenStudy (anonymous):

Yes. Sorry. I misread your explanation for c. How about (f)? I believe (d) is not a vector space.

OpenStudy (anonymous):

I agree that d is not. About f... Pretty sure it is a vector space.

OpenStudy (anonymous):

Great. Thank you.

OpenStudy (anonymous):

You're welcome.